OpenStudy (anonymous):

Lara tossed a fair coin 3 times. What is the probability of getting heads in the first two trials? 1 over 8 2 over 8 3 over 8 4 over 8 will fan and medal

terenzreignz (terenzreignz):

Okay, this might not make sense NOW, but it will later. Independent events are events that don't influence one another, and the probability of both happening is simply the product of their respective probabilities.

terenzreignz (terenzreignz):

There, I said it

terenzreignz (terenzreignz):

Now let's begin. with just one coin toss, what's the probability of getting heads? :D

OpenStudy (anonymous):

um idk

terenzreignz (terenzreignz):

Really? C'mon, at least give it a guess :)

terenzreignz (terenzreignz):

If it's easier to understand, what are the chances that you get heads with one toss?

OpenStudy (anonymous):

i honestly do not know

terenzreignz (terenzreignz):

Half. Fifty percent. 1/2 ^^

terenzreignz (terenzreignz):

Now, the second coin toss would be the same, half or 50% chance of getting heads, yes?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Are the first and second coin tosses independent of each other?

OpenStudy (anonymous):

no..?

OpenStudy (rational):

coins dont have memory

terenzreignz (terenzreignz):

in other words... does the result of the second coin toss depend on the result of the first one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

idk

OpenStudy (rational):

as if the coin remembers what happened in the previous toss and change its mind to flip the other side in next toss that can never happen

OpenStudy (anonymous):

can i just get the answer

OpenStudy (rational):

if you just blindly guess the answer, you will get it correct with a probability of 1/4 because there are 4 options

OpenStudy (anonymous):

is the answer 2/8 ?

terenzreignz (terenzreignz):

Which is a coincidence XD

OpenStudy (anonymous):

so 2/8 is the answer?

OpenStudy (rational):

Haha your luck is at its best today!

OpenStudy (anonymous):

yay lol

OpenStudy (anonymous):

can you help me with some more problems

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{2}{8}?\]

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\] which is in fact \(\frac{2}{8}\) but not what you want as a final answer

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so does anybody know the answer

OpenStudy (rational):

2/8 is correct

OpenStudy (anonymous):

At a game booth, a student gets a box of candy as the prize for winning a game. The boxes come in four colors: white, red, green, and blue. There are 10 boxes of each color. All the boxes are equally likely to be given away as prizes. Which expression shows the probability of the first winner receiving a white box and the second winner also receiving a box of the same color? 10 over 40 multiplied by 9 over 39 10 over 40 multiplied by 10 over 39 10 over 40 plus 9 over 39 10 over 40 plus 10 over 39

OpenStudy (rational):

Alright, read the question and see if you can tell me what we need to find out exactly

OpenStudy (anonymous):

2/8 is silly

OpenStudy (rational):

Haha not more sillier than 200/800 I think the teacher wants the kids to solve it by listing out all the possibilities and pick the favorable ones

OpenStudy (rational):

rational numbers trip all kids in the start because they have infinitely many representations : (2, 8) = (1, 4) = (4, 16) = ...

OpenStudy (ikram002p):

thats why i hate them :)

terenzreignz (terenzreignz):

hehe forgetting all sentimental attachment to certain box colours, the colour of the candy boxes given away to separate winners don't affect each other, right? IE... they're indpendent? ;)

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Okay, so this is a bit of a stretch, so I'm going to lay it out for you. In general, probability is: number of desirable outcomes divided by total number of possible outcomes. With forty boxes and ten white boxes, what's the probability of getting a white box?

OpenStudy (anonymous):

idk

terenzreignz (terenzreignz):

\[\Large \frac{\text{number of desired outcomes}}{\text{total number of outcomes}}\] how about now?

OpenStudy (anonymous):

still dont know

terenzreignz (terenzreignz):

Give it your best guess? I bet the choices give you some kind of clue ^^

OpenStudy (anonymous):

Why don't we try process elimination? C and D seem silly.

terenzreignz (terenzreignz):

lol yes, let's :D My professors were evil with these multiple choice questions, having the correct answer and the deceptively similar common mistakes among the choices o.O

OpenStudy (anonymous):

Lmao. ^-^

OpenStudy (anonymous):

@Skielerlucas04 are you still mad at me